A plate 60 mm wide and 12.5 mm thick is to be welded to another plate by means of parallel fillet welds. The plates are subjected to a load of 80 kN. Find the length of the weld. Assume allowable shear strength to be 60 MPa.

1
Expert's answer
2022-04-04T17:10:10-0400


Thickness of the plate to be joined

h = 12.5 mm, "\tau"allow = 60 MPa

"\tau"allow = F/A = F / (2 × 0.707h × l)

l = (8 × 104) / (2 × 0.707 × 12.5 × 60) = 75.4 mm